Tuesday, 14 June 2011

Updates 5:30pm 14 June

Current status of works is as follows:

- In addition to the earlier list of servers that have now been restored, shared web servers Storm, Volcano, and Whirlwind have also been restored this afternoon
- Shared web servers Hurricane, Drought, Blizzard and Cyclone are still being worked on by engineers. No ETA is being provided on these at this stage.
- Engineers are still working on some routing issues for some VDS servers
- Domain Registration systems are still being reviewed by engineers.  We do not expect these to be available until tomorrow at the earliest.

Please note that phone support is affected by this issue and will not be available until further notice.  We apologize for the inability to provide phone support, but given the breadth of the incident, our call centre simply cannot handle the volume of calls presently.  Please note that most staff are working outside of the office, so we are unable to provide any face-to-face support either - As we are sure you can appreciate, our focus is on getting client services operational as soon as possible, so have placed all available resources into assisting with restoration of services.